OpenStudy (anonymous):

Proof by induction. Summation from 1 to n of (3i^2-i) = n^2(n+1). Thanks for any help.

OpenStudy (anonymous):

Okay, can you show the base case is true?

OpenStudy (anonymous):

Yes. p(1); 2 = 2

OpenStudy (anonymous):

What happens when you plug \(n+1\) into the summation?

OpenStudy (anonymous):

I got \[(n+1)^{2}(n+2)\]

OpenStudy (anonymous):

No, put it into the summation.

OpenStudy (anonymous):

You should get: \[ \sum_{i=1}^{n+1}3i^2-i \]

OpenStudy (anonymous):

RIght. What you just posted is = to what I said. Is that right? Assuming we can prove it

OpenStudy (anonymous):

We can't assume that we have proved it.

OpenStudy (anonymous):

\[ \sum_{i=1}^{n+1}3i^2-i = 3(n+1)^2-(n+1) +\sum_{i=1}^{n}3i^2-i \]

OpenStudy (anonymous):

\[ =3(n+1)^2-(n+1)+n^2(n+1) \]

OpenStudy (anonymous):

Sorry for the confusion, but that is what I have so far on my paper. I am having trouble with the algebra to prove that what you just posted is equal to \[(n+1)^{2}(n+2)\]

OpenStudy (anonymous):

Excuse me, not equal to, but that your equation can be simplified down to the one I just posted.

OpenStudy (anonymous):

What does it expand to?

OpenStudy (anonymous):

Yours or mine?

OpenStudy (anonymous):

What I wrote.

OpenStudy (anonymous):

\[3n ^{2}+6n+3-n-1+n ^{3}+n^2\]

OpenStudy (anonymous):

Now expand out what it is supposed to become, and you should be able to see how you factor what you currently have into it.

OpenStudy (anonymous):

Ok. What I am trying to do is factor in such a way is that my previous post becomes \[(n+1)^2(n+2)\]

OpenStudy (anonymous):

Are we on the same page? Just making sure.

OpenStudy (anonymous):

Expand out your target and that will let you know how to factor it.

OpenStudy (anonymous):

Expanded, it is \[n^2+2n+1+n+2\]

OpenStudy (anonymous):

Where to go from here?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3(n%2B1)%5E2%E2%88%92(n%2B1)%2Bn%5E2(n%2B1) http://www.wolframalpha.com/input/?i=(n%2B1)%5E2(n%2B2) Both expand out to \(2+5 n+4 n^2+n^3\). Do you want me to teach you algebra?

OpenStudy (anonymous):

I understand that. What I am trying to do right now is simplify\[3(n+1)^2−(n+1)+n^2(n+1) \] down to \[(n+1)^2(n+2)\] I don't understand how what you just posted will help me get there.

OpenStudy (anonymous):

Simplify both expression algebraically

OpenStudy (anonymous):

Now when you simplify \((n+1)^2(n+2)\) you will see how to factor it.

OpenStudy (anonymous):

Ok. Thank you.